OpenStudy (anonymous):

Find the exact value of sec(5pi/ 3) and and sin(7pi/6)

OpenStudy (anonymous):

you know about coterminal angles?

OpenStudy (anonymous):

yeah but im not sure how to find teh coterminal angles without converting it back to degrees. that is how my teacher likes it.

OpenStudy (anonymous):

well, just realize that pi=180 degrees

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then what to i do?

OpenStudy (anonymous):

so sec x = 3/sqrt 2

OpenStudy (anonymous):

wait isnt sec 1/cos 30 = 1/ (sqrt 3)/ 2

OpenStudy (anonymous):

when i type it in teh calc it says 2 but i dont get taht when i work out teh problem

OpenStudy (anonymous):

yes you mean sqrt(3)/2 because thats a 1,sqrt(3),2 triangle and yes it is 30 degrees

OpenStudy (anonymous):

5pi/3 is 300 degrees sec(300) 1/cos(300) oh, then coterminal angle to 300 degrees is 60 degrees. so we have 1/cos(60) which is 1/1/2, which in turn is 2/1. Thus the value of sec(5pi/3) is 2

OpenStudy (anonymous):

is your calculator in radians or degrees

OpenStudy (anonymous):

you just got schooled kid

OpenStudy (anonymous):

type in your calc. the cos^-1(sqrt(3)/2))=30 degrees

OpenStudy (anonymous):

degrees

OpenStudy (anonymous):

oh i read it wrong and yes i did get schooled BIG TIME

OpenStudy (anonymous):

lol...its okay..thanks everyone for teh help...do u guys mind helping me w/ anotehr problem

OpenStudy (anonymous):

now for sin(7pi/6) 7pi/6 is 210 Which means that we have sin(210), thus the coterminal angle is 30 degree, so in reality we have sin(30), this is a know value it is 1/2. Thus sin(7pi/6) is 1/2. Thats right BIG TIME

OpenStudy (anonymous):

i don't mind so long as you realize that there is no point in converting from degrees to radians!

OpenStudy (anonymous):

or -1/2

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

trig has nothing to do with converting from some units to some other units. in fact this is a really bad idea because as functions trig functions are functions of numbers, not of angles

OpenStudy (anonymous):

yeah i was trying to figure it out w/out converting degrees to radians.

OpenStudy (anonymous):

i wasnt sure how though

OpenStudy (anonymous):

there is no point in it whatsoever

OpenStudy (anonymous):

if you see \[f(x)=x^2\] you do not ask if x is in degrees or radians or Fahrenheit or Celsius or drachmas or pound sterling

OpenStudy (anonymous):

x represents a number. and if you see \[\sin(\frac{\pi}{2})\] it means the number \[\frac{\pi}{2}\]

OpenStudy (anonymous):

watch out the math police has come to arrest us for converting radians to degrees

OpenStudy (anonymous):

damn right! your all busted. find \[\frac{7\pi}{6}\] on the unit circle on the last page of the cheat sheet. the second coordinate is sine

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@lagrange what is the derivative of sine? this is not a trick question

OpenStudy (anonymous):

in your calculator if your in radians and your typing cos(30) will NOT be the same as cos(30) in degree MOde THATS what i am trying to see what he is in.

OpenStudy (anonymous):

r u serious?

OpenStudy (anonymous):

yes perfectly

OpenStudy (anonymous):

cos (x)

OpenStudy (anonymous):

simple point i am trying to make is that the derivative of sine is NOT cosine, not if you are viewing these as functions of "degrees"

OpenStudy (anonymous):

that is as functions of numbers (not angles measured in whatever units you choose) trig functions correspond to the angle measure ONLY IN RADIANS

OpenStudy (anonymous):

yeah im serious, go get a Calculator and put them in two different modes and type it. If you dont know what mode your in than youll get the wrong answer, do you see where im coming from. Ive seen it in my trig class 100 times

OpenStudy (anonymous):

(just thought i would mention it)

OpenStudy (anonymous):

taht is exactly what my teacher says except unforutnaltey in my head i am always converting it back and forth from radians to degrees.

OpenStudy (anonymous):

so can u help me w/ anotehr problem? i wont convert it to degrees anymore.

OpenStudy (anonymous):

i understand what you are saying from the get go, but he said when he was typing it in the calc he was getting something different. (Just thought i would mention that)

OpenStudy (anonymous):

its find to convert radians to degrees and vice versa, just long as you dont get confused doing so

OpenStudy (anonymous):

just know that eventually you will have to work with radians all the time, so get used to it

OpenStudy (anonymous):

what the next problem??

OpenStudy (anonymous):

word. didnt mean to come at you, sometimes i see these chats get all heated.

OpenStudy (anonymous):

Find all teh values of x such that sin2x = sinx and \[0\le x \le 2\pi\]

OpenStudy (anonymous):

first of all subtract sinx from both sides: sin2x-sinx=0 now use the trig identity: sin2x=2sinxcosx 2sinxcosx-sinx=0 factor out sinx sinx(2cosx-1)=0 can you handle it from there?

OpenStudy (anonymous):

you have to set each factor to zero

OpenStudy (anonymous):

since sin(2x)=2sin(x)cos(x)=sin(x) 2cos(x)=1 cos(x)=1/2 since xE[0,2pi] x=pi/3 x=pi+pi/3=4pi/3

OpenStudy (anonymous):

do u find out what values of sin and cos equal 0?

OpenStudy (anonymous):

sinx(2cosx-1)=0 sinx=0 2cosx-1=0

OpenStudy (anonymous):

how do u find the other angles from 0 to 2pi

OpenStudy (anonymous):

you have to find out at what values of x doe witthin that interval is sinx =0 and cosx=1/2

OpenStudy (anonymous):

im not sure how.

OpenStudy (anonymous):

and x=pi and x=2pi is also an answer. because in the top 2sin(x)cos(x)=sin(x) doesnt matter what cos(x)= as long as the right side is the same so as long as sin(x)=0 you also satisfy this equation x=pi x=2pi

OpenStudy (anonymous):

at what values of x are sin=0 and cosx=1/2 within that interval

OpenStudy (anonymous):

if it help the answers are: 0, pi/3,pi,5pi/3, and 2pi

OpenStudy (anonymous):

when does sinx=0?? when does cosx=1/2??

OpenStudy (anonymous):

i got the 0 and pi/3. how did u get teh other asnwers. @lagra...

OpenStudy (anonymous):

look at the unit circle @ satellite provided you

OpenStudy (anonymous):

i see it now. thanks

OpenStudy (anonymous):

disreguard this pic wrong one, NOT 4PI/3 im sorry, but sin(x)=0 at pi as well as 2pi

OpenStudy (anonymous):

its 2pi-pi//3=6pi-pi/3=5pi/3

OpenStudy (anonymous):